OpenStudy (lilt):

solve 3x+3y=10, and -9x-9y=-30

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (texaschic101):

well...let me just do elimination because in this problem, it would be easier. 3x + 3y = 10 --->(3) -9x - 9y = -30 -------------- 9x + 9y = 30 (result of multiplying by 3 -9x -9y = -30 -------------add 0 = 0 This problem has INFINITE SOLUTIONS...that means whatever number you put in for x, it will work

OpenStudy (znappydooz):

Do substitution, I guess it would be easier and all